5x-10y=-5
-10x-10y=10
solve by the substitution

Answer by oberobic(2304)

5x -10y = -5
-10x -10y = 10
.
subtract the second equation from the first
.
15x = -15
x = -1
.
substituting to find y
.
5(-1) - 10y = -5
-5 -10y = -5
adding 5 to both sides
-10y = 0
so...
y = 0
.
Answer:
x = -1
y = 0
.
Done